Grade 12Algebra

Find
30.94 ÷ 0.7
Find
30.94 ÷ 0.7
Find
30.94 ÷ 0.7
Find
30.94 ÷ 0.7

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (3094/100) ÷ (7/10)
= (3094/100) × (10/7)
= (3094 × 10)/ (100 × 7)
= (442 × 1)/ (10 × 1)
= 442/10
= 44.2

Thanks